Operation Summary for July 12, 2010 - July 14, 2010 |
Monday July 12 The day shift began with Operations monitoring the Tevatron (TeV) store 7971 (107.46E30) and the Recycler stash (121.1E10), with the Antiproton Source (Pbar) stacking (22.1E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the MiniBooNE and Meson FTBF T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, the NuMI shutdown began earlier in the morning. TeV personnel started end-of-store studies (EOS) at 3:30 PM. Preaccelerator personnel switched to the H- Source. The Recycler stashed until it reached the 150E10 stash limit. During the evening shift, the TeV EOS studies finished up at 5:05 PM. Operations terminated store 7971 at 5:10 PM, put the TeV into HEP shot setup, and established store 7972 at 6:30 PM with an average initial luminosity (AIL) of 123E30. Operations had disabled the T-977 beam at 5:15 PM due to shot setup. One antiproton transfer from the Accumulator to Recycler was lost when a Recycler kicker (R:KPS2A) failed to fire. Operators searched and secured MI-31 and brought the Pelletron back online at 9:05 PM; it was operational by 10:20 PM. The Pelletron was tripping off repeatedly at the end of the evening shift and operators consulted with Electron Cooling (ECool) personnel. A MI-30 quadrupole power supply tripped off on a ground fault indication. Tuesday July 13 The midnight shift began with Operations monitoring store 7972 (55.94E30) and the stash (145.6E10), with Pbar stacking (39.7E10), and with MiniBooNE taking beam. During the midnight shift, ECool personnel came in and tuned the trajectory of the electron beam to reduce the trips. Operations terminated store 7972 at 4 AM, put the TeV into shot setup, and established store 7973 at 5:20 AM with an AIL of 236E30. MTA began its beam studies at 6:30 AM. Operations disabled the Linac beam at 7 AM and prepared the area for access; MTA personnel reported that they found their problem. Operations turned off the Booster at 8 AM for a scheduled access. During the day shift, Linac and Booster personnel came out of access at 11:30 AM. Operations reported a problem with a Linac RF station (LRF1). Operations established nominal beam intensities to MiniBooNE, T-977, and Pbar by 2 PM. During the evening shift, Operations disabled beam to T-977 at 6 PM. Operations reported a P1 line vacuum burst; operators recovered the vacuum and restored beam to Pbar. Wednesday July 14 The midnight shift began with Operations monitoring store 7973 (43.01E30) and the stash (289.5E10), with Pbar stacking (21.9E10), and with MiniBooNE taking beam. During the midnight shift, Operations terminated store 7973 at 2 AM. A TeV quadrupole corrector regulator (QB15) developed a fuse fault; operators replaced the chassis. Operations put the TeV into shot setup and established store 7974 at 4:20 AM with an AIL of 291E30. Operations established beam to T-977 at 4:45 AM. The Plans for Wednesday The plans for today are to stack, stash, and store. The Run Coordinator will allow a Pbar study, a MI coalescing and orbit study, and a D0 access later in the day. |
|
More Information Comments and Suggestions |
| last modified 07/14/2010 email Fermilab |
| Security, Privacy, Legal | |